OpenStudy (the-simple-diamond):

Macon Steinberg purchased a motor scooter with an installment loan that has an APR of 16 percent. The motor scooter sells for $1,687. The store financing requires a 20 percent down payment and 30 monthly payments. What is the finance charge? Is it $312.13??

OpenStudy (the-simple-diamond):

X=105.4375

OpenStudy (the-simple-diamond):

Where did the 2108.75 come from?

OpenStudy (the-simple-diamond):

105.4375 *4 ??

OpenStudy (the-simple-diamond):

I dont understand what your saying

OpenStudy (the-simple-diamond):

Im looking for the finance charge which= total amount repaid- amount financed Amount financed=1687

OpenStudy (the-simple-diamond):

You mean the monthly payment? 20% is the down payment which is 337.4

OpenStudy (the-simple-diamond):

Monthly payment=(1687/100) * 4.07 I got the 4.07 from a chart in my book

OpenStudy (the-simple-diamond):

Monthly payment =68.6609

OpenStudy (the-simple-diamond):

Maby we start from the top?

OpenStudy (the-simple-diamond):

https://the-simple-diamond-bunny.tumblr.com/post/154519565898 This is a chart that goes with the question

OpenStudy (the-simple-diamond):

Monthly payment=(amount of loan/100)*monthly payment for a $100 loan Total amount repaid= number of payments* monthly payments Finance charge = total amount repaid-amount financed

OpenStudy (the-simple-diamond):

Monthly payment=(1687/100)*4.07=68.6609

OpenStudy (the-simple-diamond):

Total amount repaid=30*68.6609=2059.827

OpenStudy (the-simple-diamond):

Finance charge=2059.827-1687=372.827

OpenStudy (the-simple-diamond):

Ive triple cheked my math and i get the same answer but i found the exact same wuestion on a multiple choice quiz online A. $300.98 B. $298.90 C. 312.13 D. 325.00 So i dont know what im doing wrong

OpenStudy (mrs.ambrose614):

It's c

OpenStudy (the-simple-diamond):

Thanks